Gotham Bridge and Tunnel Authority practices peak-load pricing. During peak hours when bridges are at capacity...

  1. Gotham Bridge and Tunnel Authority practices peak-load pricing. During peak hours when bridges are at capacity and the price elasticity of demand is −3, Gotham charges $25 per bridge crossings. During off-peak hours the price elasticity of demand is −5. How much should Gotham charge during off-peak hours to maximize profit?

  • $18

  • $24

  • $20

  • $21

  • There is not enough information to answer this question.
Answer #1

Correct Answer:

D

Working note:

As per the Lerner index,

(P-MC)/P = 1/|E|

(25 - MC)/25 = 1/3

MC = 25 - 25/3

MC = $16.67

When E = -5

(P - 16.67)/P = 1/5

5P - 5*16.67 = P

P = 5*16.67/4

P = $20.84 or $21

So, price should be $21.